### v3

Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, charts that distort the underlying data, leading readers to draw inaccurate conclusions that may support disinformation. Here, we uncover an important vulnerability: MLLM question-answering accuracy on misleading visualizations drops on average to the level of a random baseline. To address this, we introduce the first inference-time methods to improve performance on misleading visualizations, without compromising accuracy on non-misleading ones. The most effective method extracts the underlying data table and uses a text-only LLM to answer the question based on the table. Our findings expose a critical blind spot in current research and establish benchmark results to guide future efforts in reliable MLLMs. [^1] Code and data available at: github.com/UKPLab/arxiv2025-misleading-visualizations
keywords:
large language models, chart understanding, visualization
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, or companies to efficiently communicate data insights to a broad audience [1]. They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public about the spread of the disease [2]. In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers [3]. However, visualizations can also be misleading [4]. This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader [5, 6, 4, 7, 8, 9, 10, 11, 12]. These misleaders appear across a wide range of visualizations, including bar, pie, and line charts, choropleth maps, or scatter plots. Recent taxonomies have documented over 70 distinct types of misleaders observed in real-world examples [7, 9]. Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like area encoding and 3D effects. Figure 1 shows three real-world examples of misleading visualizations [7] paired with multiple-choice questions. Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer (in purple) instead of the correct one (in green). Figure 2 illustrates how two visualizations of the same data, one non-misleading and the other misleading, can lead to different interpretations by readers.
Figure 1: Three examples of real-world misleading visualizations [7] with multiple choice questions. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
Figure 2: A pair of bar charts, one non-misleading and one misleading, representing the same underlying data [11]. Both are paired with the same Likert-scale question, where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses for both visualizations. Assuming a score of 3 is given for the non-misleading chart, only a score of 3 is valid for the misleading one. However, the truncated axis deceives readers in choosing a value higher than 3.
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies [6, 13, 14, 15] have shown that such visualizations reduce human accuracy when answering multiple-choice, free-text, or rank questions about the underlying data. Other studies [6, 16, 11] have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
Misleading visualizations pose a serious threat to society. Due to their high deceptive potential, they can be exploited by malicious actors to promote online disinformation [4]. For instance, the visualizations in Figure 1 mislead readers on sensitive topics such as COVID-19, abortion, and gun violence statistics. Misleading visualizations can shift public opinion, even on highly polarized political debates like Brexit [17].
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research [1], with steady progress on reference benchmarks such as ChartQA [18, 19], highlighting their strong potential as interactive assistants for extracting data insights. However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it. These risks underscore the urgency of thoroughly assessing the vulnerability of MLLMs to misleading visualizations and developing effective mitigation strategies.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 18 MLLMs of varying sizes to misleading visualizations. This collection includes 11 open-weight MLLMs from the Llava-Next [20], Qwen2VL [21], Ovis-1.6 [22], and InternVL2.5 [23] families, ranging from 2 to 38 billion parameters. Additionally, we include four commercial models, GPT4 [24], GPT4o [24], Gemini-1.5-flash [25], and Gemini-1.5-pro [25], as well as three open-weight MLLMs specifically trained for visualization understanding: Llava-Chart-Instruct [26], TinyChart [27], and ChartGemma [28] ranging from 3 to 13 billion parameters.
In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Table 1 [13, 14, 15, 7]. It combines the existing synthetic CALVI ($n$ = 45) [13] and CHARTOM ($n$ = 56) [14, 15] datasets with 42 real-world examples [7] that we manually annotated with multiple choice questions; (b) a non-misleading visualization dataset ($n$ = 124) [13, 14, 15, 29], also combining real-world and synthetic cases; and (c) ChartQA [18], the reference real-world non-misleading benchmark ($n$ = 2500). In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. We use the synthetic dataset of Lauer & O’Brien, which consists of four such visualization pairs [11]. Detailed statistics about the datasets, including the number of misleader categories and the type of questions, are reported in Table 2.
Experiments 1 and 2 demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. Experiment 1 yields several key insights: QA accuracy drops by up to 65.5 percentage points (pp) compared to ChartQA, with half the MLLMs failing to outperform the random baseline; MLLMs are less vulnerable to real-world misleading visualizations than to synthetic ones; and vulnerability varies substantially depending on the type of misleader.In addition, Experiment 2 shows that MLLMs exhibit inconsistency when assigning Likert-scale scores to two visualizations of the same data when one of them has a truncated axis.
To mitigate the high vulnerability of MLLMs observed in Experiment 1, we introduce six inference-time correction methods in Experiment 3. These methods modify the input image and/or prompt without requiring additional training data or fine-tuning. We evaluate them on both the misleading and non-misleading visualization datasets used in Experiment 1. The goal is to improve QA accuracy on misleading visualizations while maintaining performance on non-misleading ones.
We compare the following methods, illustrated in Figure 3: (1) including a warning message to the prompt to make the MLLM aware of specific misleaders in the visualization; (2) extracting the axes using the MLLM and incorporating them into the prompt; (3) extracting the underlying data table and adding it into the prompt; (4) combining both extracted axes and table into the prompt; (5) providing the extracted table, without the visualization, to a text-only LLM, reframing the task as table-based QA; and (6) using a text-only LLM to generate code for a new visualization based on the extracted table, which then replaces the original visualization in the QA prompt.
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 18 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We are the first to introduce six correction methods to mitigate the negative impact of misleaders on QA performance, while maintaining accuracy on non-misleading visualizations.
The results of the three experiments expose a critical blind spot in current automated chart understanding research, highlighting the need for future work to make MLLMs more robust to misleading visualizations.
Figure 3: Illustration of the six inference-time correction methods applied on a misleading visualization from CALVI [13]. The visualization suffers from inconsistent tick intervals on the y-axis. The correct answer is True. In this case, it is only predicted after redrawing the visualization.
1 Results
1.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Experiment 1 assesses the ability of MLLMs to answer questions about the underlying data given a visualization. The upper part of Figure 4 presents the QA accuracy of 18 MLLMs across the three datasets, sorted by increasing accuracy on ChartQA. The bottom part displays the accuracy for the three subsets constituting the misleading visualizations dataset. The results of Experiment 1 reveal three key findings.
Figure 4: Top: Accuracy (%) of various MLLMs on misleading visualization [13, 14, 15, 7], non-misleading visualization [13, 14, 15, 29], and ChartQA datasets [18]. The horizontal dashed line indicates the accuracy of the random baseline on misleading visualizations. Models are sorted by increasing accuracy on ChartQA. Bottom: Accuracy (%) of various MLLMs on subsets of the misleading visualizations. The horizontal dashed lines indicate average human accuracy on CALVI and CHARTOM [13, 14, 15].
1.1.1 Performance is worse on misleading visualizations
All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 pp, with an average of 23.4 pp. The decline is even more pronounced when compared to ChartQA, reaching up to 65.5 pp, with an average of 54.0 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (25.6%) is on par with the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret the distorted data, making them highly vulnerable to misleaders.
For comparison, average human accuracy reported in prior work on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively [13, 14, 15]. On the same datasets, MLLMs achieve an average of 51 and 47% on non-misleading visualizations and 24 and 21% on misleading ones, falling short of human performance in all cases.
The ranking of MLLMs by accuracy varies a lot across the three datasets. Notably, the two best-performing models on the CALVI subset of misleading visualizations rank among the worst on ChartQA, placing second and third from the bottom. This discrepancy suggests that achieving high accuracy on misleading visualizations is unlikely to emerge naturally from improvements on ChartQA, underscoring the need for dedicated mitigation methods.
1.1.2 Performance is higher on real-world misleading visualizations
MLLMs perform better on the real-world subset of the misleading visualizations dataset, achieving an average accuracy of 33.5%, compared to 24% on CALVI and 21% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, GPT4o, Gemini-1.5-pro, and InternVL2.5-38B, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events and statistics, without relying on the content of the visualization. In such cases, parametric knowledge effectively serves as a form of protection against real-world misleaders. However, this finding should be interpreted with caution. In practical scenarios, the underlying data will often be recent and therefore unlikely to be covered by the model’s parametric knowledge.
1.2 Performance varies by type of misleader
Figure 5: Average Likert-scale ratings (1 to 6) of MLLMs and human readers for the four pairs of misleading and non-misleading visualizations depicting the same underlying data from the dataset of Lauer & O’Brien [11]. MLLM results are reported with standard deviations. Human results are reported from prior work [11].
Not all misleading visualizations are equally deceptive. This was already observed with human readers [13, 14, 15], and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Figure 7. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 64.8, 50.0, and 43.7, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and area encoding and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 23.5 and 27.4 %, respectively.
There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader [14]. Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects [13]. The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM [14], compared to MLLMs, which achieve an average of 35.4%. These results highlight a notable difference in how humans and MLLMs process visualizations and are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers [30].
1.3 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Experiment 2 evaluates whether MLLMs interpret consistently two visualizations that represent the same underlying data table, one misleading and the other non-misleading. This is assessed using a Likert-scale question, such as in Figure 2. Unlike in Experiment 1, there is no ground truth answer. Instead, since both visualizations depict the same underlying data, a model is considered misled if it assigns different scores to each of them. We assess the significance of the difference in ratings with a two-sided paired t-test (p-value (p) $\leq$ 0.05).
Figure 5 summarizes the average Likert-scale ratings from all 18 MLLMs for each visualization pair. Human results from prior work [11] are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is statistically significant for the two visualization pairs involving truncated axis: one with bar charts (Test statistic (T) = -3.2, p = 5e-3) and one with line charts (T = -3.0, p = 0.01). For the other two pairs, misrepresentation with bubble charts (T = -0.77, p = 0.45) and 3D effects with pie charts (T = -0.62, p = 0.55), the differences are not statistically significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, in particular for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s answers consistency.
1.4 Experiment 3 - Correction methods
Having established that MLLMs are vulnerable to misleading visualizations, we explore six correction methods aimed at improving their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline on misleading visualizations in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) [31]. All test statistics and p-values are reported in Table 3. Figure 6 shows the change in accuracy relative to Experiment 1 across the six methods on the misleading and non-misleading visualization datasets: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization. Qwen2.5-7B [32] is used as the text-only LLM for both table-based QA and redrawing visualizations.
Figure 6: Change ($\Delta$) in accuracy (percentage points) compared to Experiment 1 on misleading and non-misleading visualization datasets using different inference-time correction methods to mitigate misleading visualizations. Statistically significant changes ($\leq$0.05) are hashed.
1.4.1 Two methods are effective: table-based QA and redrawing the visualization
The most effective approach by far to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Its strength lies in fully replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis, or misleading annotations. However, this method proves less effective against misleaders that persist even when the data is presented in table form, such as inappropriate item order and concealed uncertainty. Table-based QA comes with a notable cost for non-misleading visualizations. Two models experience significant decreases in accuracy, by up to 8.5 pp. This degradation is primarily due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost.
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations [30]. It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib [33], which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%), and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer and more detailed Python code. For all other chart types, the redrawing success rate exceeds 90%. Unlike table-based QA, this method does not result in significant decreases in accuracy on non-misleading visualizations, making it a more attractive option than table-based QA when preserving performance on non-misleading visualization is a priority.
1.4.2 Table extraction is a key intermediate step
Both table-based QA and redrawing the visualization depend heavily on the accuracy of an intermediate step: table extraction. Since the underlying data table is not available, it must be derived from the visualization before either correction method can be applied. This extraction step is often non-trivial, particularly for complex chart types such as maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. Perfect accuracy in table extraction is not always necessary, its importance depends on the nature of the question. Some questions can be answered correctly as long as the extracted table preserves the overall trend of the underlying data, while others require exact values. As discussed above, errors in table extraction can significantly decrease accuracy on non-misleading visualizations. In the absence of ground truth tables, we cannot directly evaluate the accuracy of table extraction. Instead, we analyze MLLM outputs on CHARTOM [14, 15], which pairs each question with two visualizations, one misleading and one non-misleading, generated from the same underlying data. Ideally, the tables extracted from both visualizations should be identical. However, for all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). Figure 8 illustrates the extraction of two different tables from a CHARTOM pair. Given that the two most effective correction methods depend on reliable table extraction, and that extraction errors also negatively impact performance on non-misleading visualizations, improving this intermediate step is a key direction for future work.
1.4.3 Other correction methods are not effective
None of the remaining correction methods yield significant improvements. Including the extracted table in the prompt even results in a significant decrease in accuracy on non-misleading visualizations for one MLLM. While using the extracted table alone proves to be the most effective method, combining it with the original misleading visualization does not produce similar gains. This further suggests that removing or replacing the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method. Although many misleaders rely on axis manipulations, providing the extracted axes alone has little to no effect on performance. Similarly, adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task [34, 35]. Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing to select the most suitable correction method based on the type of misleader.
2 Discussion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Experiment 1 demonstrated that MLLMs are unable to accurately interpret the underlying data when presented with misleading visualizations, reducing their QA accuracy, on average, to the level of a random baseline. Experiment 2 showed that MLLMs are particularly susceptible to truncated axes, often failing to provide consistent answers when shown misleading and non-misleading versions of the same data. Identifying this vulnerability fills a critical gap in automated chart understanding research. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation [4] in polarized political contexts [17] or during crises like the COVID-19 pandemic [36]. Addressing this blind spot is an urgent priority for future research in automated chart understanding.
To mitigate this issue, we proposed six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, table-based QA comes at the cost of reduced accuracy on non-misleading visualizations. While most recent work in chart understanding treats visualizations purely as images [26, 27, 28], our findings challenge this direction and sheds a new light on alternative approaches that involve table extraction as an intermediate step [37, 38, 39].
Our correction methods, in particular redrawing the visualization, relate to prior work on visualization linters [40, 41, 30]: rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations [30], they can only operate with specific chart design tools such as Matplotlib [33, 40, 41] or a very limited set of chart and misleader types [30]. In contrast, our correction methods do not impose such restrictions.
Our findings are further supported by prior work [42] and parallel studies [43, 44, 45], which also report MLLM vulnerabilities to misleaders [42, 43, 44] and to simple design decisions like the color palette [45], reinforcing the urgency and relevance of this new research direction. Our study differentiates itself in several key ways. Prior work [42] only evaluated one MLLM, GPT4, on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world misleading visualizations and show that vulnerability is somewhat lower in these cases, especially for larger models which may rely on parametric knowledge. Beyond identifying the vulnerability, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works [44].
We identify three limitations to our work. First, the visualization redrawing method does not support maps, as Matplotlib [33] lacks sufficient functionality for rendering high-quality maps. Second, we assume prior knowledge of the chart type (bar, line, …) for generating the prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step. Third, existing datasets do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI [13] and present to a lower extent in the real-world cases [7].
3 Methods
3.1 QA datasets
The misleading and non-misleading visualization QA datasets used in Experiments 1 and 3 combine three existing resources and one introduced in this work.
First, CALVI [13] includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with a multiple-choice question (MCQ) with three to four choices. The dataset was initially designed to evaluate human readers.
Second, CHARTOM [14, 15] contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. The MCQs provide two to four choices. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. We changed the name of certain misleaders in CHARTOM to align them with the terminology used in CALVI [7, 13].
Third, VLAT [29], the reference dataset to evaluate the human comprehension of visualizations, provides 12 non-misleading visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. MCQs have two to four choices.
Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with a MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them [7], which inspired the synthetic examples in CALVI. We manually create MCQs, using as reference those from CALVI and CHARTOM sharing the same misleader. The motivation for creating this additional resource is that CALVI and CHARTOM rely both on synthetic data. By incorporating questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
The non-misleading visualization dataset combines the synthetic instances of CALVI ($n$ = 15) of CHARTOM ($n$ = 56) with the real-world ones of VLAT ($n$ = 53).
We also provide the previously reported performance of MLLMs on the test set of ChartQA [18], which contains 2500 real-world visualizations paired with free-text questions, of which half are human-written and the others are AI-generated.
The datasets cover together 12 chart types: line charts, area charts, stacked area charts, bar charts, stacked bar charts, pictorial bars, histograms, pie charts, scatter plots, bubble charts, maps, treemaps.
For free-text questions where the expected answer is a number, we adopt the relaxed accuracy metric of ChartQA [18]. Specifically, a prediction is considered correct if it falls within a ±5% interval of the ground truth. For MCQs, we present the choices in their default order. In Table 4, we report the mean accuracy of the MLLMs over three different choice orders. The choices are shuffled using three randomly generated numbers as seeds (654, 114, 25). Standard deviations for the accuracy range from 0.0 to 4.06 pp.
3.2 Likert-scale dataset
For Experiment 2, we use four pairs of visualizations, one misleading and the other non-misleading, originally designed to evaluate humans [11]. The four pairs cover four unique chart types and three distinct misleaders.
3.3 Inference with (M)LLMs
All prompts are available in the code, which is provided as supplementary material. We use the Hugging Face transformers [46] implementation for all open-weight (M)LLMs and access GPT and Gemini models via the Azure OpenAI API and Google Cloud API, respectively. To ensure deterministic outputs, we set the random seed at 42, set the temperature to 0, and use a top-p value of 1. Following the standard ChartQA evaluation setup, all (M)LLMs are prompted in a zero-shot manner. For TinyChart, we report results using the Direct approach [27].
3.4 Implementation of correction methods
Misleader warning: we insert in the prompt a short warning message based on the definitions of misleaders. The message is the same for all instances with the same type of misleader. There are no messages for the five types of misleaders where the visualization is deceiving only in the context of a specific question, making a standardized warning message impossible: cherry picking, misleading annotations, concealed uncertainty, missing normalization, and missing data.
Axes and table extraction: we prompt the MLLM to extract the axes or underlying table in a zero-shot setting. The axes and tables are formatted as text strings. We do not impose constraints on the delimiters used to indicate new rows and columns.
Table-based QA: the extracted table is provided as input in the prompt to the text-only LLM with the question and an instruction to answer it based on the table.
Redrawn visualization: the text-only LLM receives the extracted table and the chart type as text input and generates Python code to create a visualization using Matplotlib [33]. The code is then executed to produce a new visualization. If the code compiles successfully, the newly generated visualization replaces the original one in the QA prompt; otherwise, the original visualization remains in use.
[…]
Competing interests
The authors declare no competing interests.
[…]
Supplementary information
The datasets and code to reproduce the experiments are made available at github.com/UKPLab/arxiv2025-misleading-visualizations.
[…]
Acknowledgements
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We want to express our gratitude to Niklas Traser for conducting an initial exploration of the real-world data, to Jan Zimny for our insightful discussions on the topic of misleading visualizations, and to Germàn Ortiz, Manisha Venkat, and Max Glockner for their feedback on a draft of this work.
Appendix A Extended data - Definition of misleaders
Table 1: Extended data - The 17 types of misleader included in this work, with their number of occurrences ($n$) and definitions [7, 13].
Misleader Definition
Inverted axis ($n$=26) An axis is oriented in an unconventional direction
and the perception of the data is reversed [7].
Truncated axis ($n$=21) The axis does not start from zero or is truncated
in the middle resulting in an exaggerated difference
between the two bars [7].
Inappropriate axis range ($n$=15) The axis range is either too broad or too narrow to
accurately visualize the data, allowing changes to be
minimized or maximized depending on
the author’s intention [7].
Inconsistent tick intervals ($n$=12) Cases with varying intervals between the ticks [7].
3D effects ($n$=12) The closer something is, the larger it appears,
despite being the same size in 3D perspective [7].
Inappropriate item order ($n$=9) The axis labels or legends appear to be in a random
order due to manipulation of data ordering [13].
Inappropriate aggregation ($n$=8) Aggregating data in an improper way that leads to
inaccurate conclusions [13].
Dual axis ($n$=8) Two independent axes are layered on top of
each other with inappropriate scaling [7].
Misrepresentation ($n$=7) The value labels provided do not match
the visual encoding [7].
Cherry picking ($n$=6) Selecting only a subset of data to display, which
can be misleading if one is asked to infer something
about the whole set of data [13].
Misleading annotations ($n$=5) Annotations that contradict or make it harder
to read the visualization [13].
Area encoding ($n$=5) Linearly encoding the values as areas
leads the readers to consistently
underestimate the values [7].
Concealed uncertainty ($n$=3) Not displaying uncertainty in visualizations may
misrepresent the certainty in the underlying data.
In the case of prediction making, this can misguide
the viewers to falsely overconfident conclusions [13].
Missing normalization ($n$=3) Displaying unnormalized data in absolute quantity
when normalized data in relative quantity
is of interest [13].
Inappropriate use of pie chart ($n$=1) When a pie chart is used for non-part-to-whole data,
it creates confusion for the audience, who may
misinterpret the significance of a given section [7].
Missing data ($n$=1) A visual representation implies data exist but
the data is actually missing [13].
Overplotting ($n$=1) Displaying too many things on a plot can
obscure parts of the data [13].
[…]
Appendix B Extended data - Datasets statistics
Table 2: Extended data - Datasets statistics. $n$ is the number of instances (question-answer-visualization triplets). $m$ is the number of misleader categories. MCQ indicates multiple-choie questions.
Dataset Experiments $n$ $m$ Question types
Misleading visualizations 1 & 3 143 17 MCQ, Free-text, rank
$\hookrightarrow$ CALVI (misleading) [13] 1 & 3 45 14 MCQ
$\hookrightarrow$ CHARTOM (misleading) [14, 15] 1 & 3 56 7 MCQ, Free-text, rank
$\hookrightarrow$ Real-world [7] 1 & 3 42 12 MCQ
Non-misleading visualizations 1 & 3 124 - MCQ, Free-text, rank
$\hookrightarrow$ CALVI (non-misleading) [13] 1 & 3 15 - MCQ
$\hookrightarrow$ CHARTOM (non-misleading) [14, 15] 1 & 3 56 - MCQ, Free-text, rank
$\hookrightarrow$ VLAT [29] 1 & 3 53 - MCQ
ChartQA [18] 1 2500 - Free-text
Lauer & O’Brien [11] 2 8 3 Likert-scale
Appendix C Extended data - Experiment 1 - Detailed accuracy scores
Figure 7: Extended data - Detailed accuracy scores (%) of Experiment 1 per misleader and MLLM.
Appendix D Extended data - Experiment 3 - significance tests
Table 3: Extended data - Test statistics (T) and P-values (p) for the McNemar test assessing the significance of the correction methods in Experiment 3. Significant results ($\leq$0.05) are marked in bold.
Method Model Misleading Non-misleading
T p T p
qwen2vl 5 1.00 0 1.00
Misleader warning ovis1.6 6 0.17 1 0.63
internvl2.5 4 1.00 0 1.00
qwen2vl 7 1.0 1 0.38
Axes ovis1.6 8 0.38 8 0.21
internvl2.5 5 1.00 5 1.00
qwen2vl 8 0.15 7 1.00
Table ovis1.6 9 0.40 11 0.04
internvl2.5 9 0.82 6 0.12
qwen2vl 6 0.12 6 0.45
Table + Axes ovis1.6 11 1.00 10 0.10
internvl2.5 8 1.00 10 0.68
qwen2vl 12 1e-4 10 0.05
Table-based QA ovis1.6 16 3e-3 10 0.05
internvl2.5 12 1e-4 17 0.43
qwen2vl 5 0.01 2 0.11
Redrawing visualization ovis1.6 5 2e-4 6 0.33
internvl2.5 11 0.84 6 0.24
Appendix E Extended data - Example of table extraction
Figure 8: Extended data - Table extraction applied to a non-misleading instance of CHARTOM and its misleading version with an inverted axis. Despite that the two visualizations represent the same data, the extracted tables show different values and trends.
Appendix F Extended data - Experiment 1 - Impact of choices order in MCQs
Table 4: Extended data - Mean accuracy (%) and standard deviation when shuffling MCQ choices of Experiment 1 with three different random seeds.
Model Misleading visualizations Non-misleading visualizations
llava-7B 24.48 $\pm$ 0.70 44.35 $\pm$ 2.79
llava-13B 24.47 $\pm$ 2.42 37.36 $\pm$ 4.06
qwen2vl-2B 24.71 $\pm$ 3.85 38.44 $\pm$ 1.23
tinychart-3B 22.15 $\pm$ 1.76 22.04 $\pm$ 1.23
gpt4 30.54 $\pm$ 2.14 47.58 $\pm$ 2.91
llava-chart-instruct-13B 27.97 $\pm$ 2.52 34.14 $\pm$ 2.33
internvl2.5-2B 19.58 $\pm$ 3.21 43.28 $\pm$ 1.23
chartgemma-3B 22.61 $\pm$ 2.82 26.61 $\pm$ 2.91
ovis1.6-9B 23.55 $\pm$ 0.40 52.15 $\pm$ 2.83
qwen2vl-7B 20.75 $\pm$ 2.25 48.93 $\pm$ 3.26
internvl2.5-4B 19.35 $\pm$ 2.25 46.50 $\pm$ 2.46
internvl2.5-8B 22.61 $\pm$ 1.07 55.38 $\pm$ 1.68
ovis1.6-27B 28.44 $\pm$ 0.40 53.76 $\pm$ 1.23
gpt4o 37.76 $\pm$ 0.00 59.95 $\pm$ 1.86
internvl2.5-26B 20.05 $\pm$ 2.46 55.38 $\pm$ 1.86
internvl2.5-38B 32.64 $\pm$ 1.07 62.63 $\pm$ 0.46