### v4

# Protecting multimodal large language models against misleading visualizations

Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, i.e., charts that distort the underlying data, leading readers to draw inaccurate conclusions that may support disinformation. Here, we uncover an important vulnerability: MLLM question-answering (QA) accuracy on misleading visualizations drops on average to the level of the random baseline. To address this, we introduce the first inference-time methods to improve QA performance on misleading visualizations, without compromising accuracy on non-misleading ones. We find that two methods, table-based QA and redrawing the visualization, are effective, with improvements of up to 19.6 percentage points. We make our code and data available.111github.com/UKPLab/arxiv2025-misleading-visualizations
1 Introduction
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, or companies to efficiently communicate data insights to a broad audience (Huang et al., 2025). They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public about the spread of the virus (Zhang et al., 2021). In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers (Pandey et al., 2014). However, visualizations can also be misleading (Correll and Heer, 2017). This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Correll and Heer, 2017; Lauer and O’Brien, 2020; McNutt et al., 2020; Yang et al., 2021; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). These misleaders appear across a wide range of visualizations, including bar, pie, and line charts, choropleth maps, or scatter plots. Recent taxonomies have documented over 70 distinct types of misleaders observed in real-world examples (Lo et al., 2022; Lan and Liu, 2025). Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like area encoding and 3D effects. Figure 1 shows three real-world examples of misleading visualizations (Lo et al., 2022) paired with multiple-choice questions. Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer instead of the correct one. Figure 2 illustrates how two visualizations of the same underlying data, one non-misleading and the other misleading, can lead to different interpretations by readers.
Figure 1: Three examples of real-world misleading visualizations (Lo et al., 2022) with multiple choice questions. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
Figure 2: Non-misleading and misleading visualizations of the same data (Lauer and O’Brien, 2020) with a Likert-scale question where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses for both visualizations. However, the truncated axis deceives readers in choosing a higher value for the misleading visualization.
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies (Pandey et al., 2015; Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024) have shown that such visualizations reduce human accuracy when answering multiple-choice, free-text, or rank questions about the underlying data. Other studies (Pandey et al., 2015; O’Brien and Lauer, 2018; Lauer and O’Brien, 2020) have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
Misleading visualizations pose a serious threat to society. Due to their deceptive potential, they can be exploited by malicious actors to promote online disinformation (Correll and Heer, 2017). For instance, the charts in Figure 1 mislead readers on sensitive topics such as COVID-19, abortion, and gun violence. Misleading visualizations can shift public opinion, even on polarized debates like Brexit (Tartaglione and de Wit, 2025).
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research (Huang et al., 2025), with steady progress on reference benchmarks such as ChartQA (Masry et al., 2022; Hoque et al., 2022), highlighting their strong potential as interactive assistants for extracting data insights. However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 18 MLLMs of varying sizes to misleading visualizations. In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs on misleading and non-misleading visualizations. In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. The results of both experiments demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. In Experiment 3, we introduce six inference-time correction methods to reduce the vulnerability of MLLMs to misleading visualizations in a QA setting. We evaluate them on the datasets of Experiment 1. The goal is to improve QA accuracy on misleading visualizations while maintaining performance on non-misleading ones.
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 18 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We are the first to introduce six correction methods to mitigate the negative impact of misleaders on QA tasks.
2 Methodology
2.1 Datasets
Experiments 1 and 3 We compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Table 2 (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lo et al., 2022); (b) a non-misleading visualization dataset ($n$ = 124) (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lee et al., 2017), also combining real-world and synthetic cases; and (c) ChartQA (Masry et al., 2022), the reference real-world non-misleading benchmark ($n$ = 2500). Datasets (a) and (b) combine three existing resources and one introduced in this work. First, CALVI (Ge et al., 2023) includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with a multiple-choice question (MCQ). The dataset was initially designed to evaluate humans. Second, CHARTOM (Rho et al., 2023; Bharti et al., 2024) contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. Third, VLAT (Lee et al., 2017), the reference dataset to evaluate humans, provides 12 non-misleading visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with a MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them (Lo et al., 2022), which inspired the synthetic examples in CALVI. We manually create MCQs, using as reference those from CALVI and CHARTOM. By incorporating visualizations and questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
Experiment 2 We use four pairs of visualizations, one misleading and the other non-misleading, originally designed by Lauer and O’Brien (2020) to evaluate humans. The pairs cover three distinct misleaders.
Table 1 provides detailed dataset statistics.
Dataset Experiments $n$ $m$ Question types
Misleading visualizations 1 & 3 143 17 MCQ, Free-text, rank
$\hookrightarrow$ CALVI (misleading) 1 & 3 45 14 MCQ
$\hookrightarrow$ CHARTOM (misleading) 1 & 3 56 7 MCQ, Free-text, rank
$\hookrightarrow$ Real-world 1 & 3 42 12 MCQ
Non-misleading visualizations 1 & 3 124 - MCQ, Free-text, rank
$\hookrightarrow$ CALVI (non-misleading) 1 & 3 15 - MCQ
$\hookrightarrow$ CHARTOM (non-misleading) 1 & 3 56 - MCQ, Free-text, rank
$\hookrightarrow$ VLAT 1 & 3 53 - MCQ
ChartQA 1 2500 - Free-text
Lauer & O’Brien 2 8 3 Likert-scale
Table 1: Datasets statistics. $n$ is the number of instances. $m$ is the number of misleader categories.
2.2 Correction methods
Figure 3: Illustration of the six inference-time correction methods applied to a misleading visualization from CALVI (Ge et al., 2023). The visualization suffers from inconsistent tick intervals on the y-axis.
In Experiment 3, we compare six inference-time correction methods, illustrated in Figure 3, to increase QA accuracy on misleading visualization without compromising it on non-misleading ones.
(1) Misleader warning: we insert in the prompt a message to make the MLLM aware of the misleader in the visualization. The message is identical for all instances with the same misleader.
Providing (2) the axes data, (3) the data table, (4) or both: we prompt the MLLM in zero-shot to extract the axes labels or underlying data table. The axes and tables are formatted as text strings. We do not impose constraints on the delimiters used to indicate new rows and columns. The axes, the table, or both are then provided as additional input in the prompt.
(5) Table-based QA: after extracting the table with the MLLM, we provide it alone to a text-only LLM, reframing the task as table-based QA.
(6) Redrawn visualization: after extracting the table with the MLLM, we provide it to a text-only LLM which generates Python code to create a visualization using Matplotlib (Hunter, 2007). If the code compiles successfully, the generated visualization replaces the original one in the QA prompt; otherwise, the original one is used.
2.3 Implementation details
We conduct experiments with 18 MLLMs, including 11 open-weight MLLMs from the Llava-Next (Liu et al., 2024), Qwen2VL (Wang et al., 2024), Ovis-1.6 (Lu et al., 2024), and InternVL2.5 (Chen et al., 2024) families, ranging from 2 to 38 billion parameters. Additionally, we include four commercial models, GPT4 (OpenAI, 2023), GPT4o (OpenAI, 2023), Gemini-1.5-flash (Gemini-Team, 2024), and Gemini-1.5-pro (Gemini-Team, 2024), as well as three open-weight MLLMs specifically trained for visualization understanding: Llava-Chart-Instruct (Zeng et al., 2025), TinyChart (Zhang et al., 2024), and ChartGemma (Masry et al., 2025) ranging from 3 to 13 billion parameters. Qwen2.5-7B (Qwen-Team, 2024) serves as the text-only LLM. We use the transformers library (Wolf et al., 2020) to access all open models and the Azure API and Google Cloud API for GPT and Gemini models, respectively. We set the random seed at 42 and the temperature to 0. Following the standard ChartQA setup, all (M)LLMs are prompted in zero-shot. We report TinyChart results with the Direct approach (Zhang et al., 2024).
3 Results
3.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Figure 4: Top: Accuracy (%) of various MLLMs on misleading visualization, non-misleading visualization, and ChartQA datasets. The horizontal dashed line indicates the accuracy of the random baseline on misleading visualizations. Models are sorted by increasing accuracy on ChartQA. Bottom: Accuracy (%) of various MLLMs on subsets of the misleading visualizations. The horizontal dashed lines indicate average human accuracy on CALVI and CHARTOM (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024).
Experiment 1 assesses the ability of MLLMs to answer questions about the underlying data given a visualization. The upper part of Figure 4 presents the QA accuracy of 18 MLLMs across the three datasets, sorted by increasing accuracy on ChartQA. The bottom part displays the accuracy for the three subsets constituting the misleading visualizations dataset. The results of Experiment 1 reveal three key findings.
Performance is worse on misleading visualizations. All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 percentage points (pp), with an average of 23.4 pp. The decline is even more pronounced compared to ChartQA, reaching up to 65.5 pp, with an average of 54.0 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (25.6%) is on par with the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret data distorded by misleaders.
For comparison, average human accuracy reported on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024). On the same datasets, MLLMs achieve an average of 51 and 47% on non-misleading visualizations and 24 and 21% on misleading ones, falling short of human performance in all cases.
The ranking of MLLMs by accuracy varies a lot across the three datasets. Notably, the two best-performing models on the CALVI subset of misleading visualizations rank among the worst on ChartQA, placing second and third from the bottom. This discrepancy suggests that achieving high accuracy on misleading visualizations is unlikely to emerge naturally from improvements on ChartQA, underscoring the need for dedicated mitigation methods.
Performance is higher on real-world misleading visualizations than on other subsets. MLLMs perform better on the real-world subset of the misleading visualizations dataset, achieving an average accuracy of 33.5%, compared to 24% on CALVI and 21% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, GPT4o, Gemini-1.5-pro, and InternVL2.5-38B, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events and statistics, without relying on the content of the visualization. In such cases, parametric knowledge effectively serves as a form of protection against real-world misleaders. However, this finding should be interpreted with caution. In practical scenarios, the underlying data will often be recent and therefore unlikely to be covered by the model’s parametric knowledge.
Figure 5: QA Accuracy (%) per type of misleader.
Performance varies by type of misleader. Not all misleading visualizations are equally deceptive. This was already observed with human readers (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024), and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Figure 5. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 64.8, 50.0, and 43.7, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and area encoding and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 23.5 and 27.4 %, respectively. There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader (Rho et al., 2023). Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects (Ge et al., 2023). The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM (Rho et al., 2023), compared to MLLMs, which achieve an average of 35.4%. These results highlight a notable difference in how humans and MLLMs process visualizations and are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers (Fan et al., 2022).
3.2 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Figure 6: Average Likert-scale ratings (1 to 6) on four pairs of misleading and non-misleading visualizations depicting the same data. MLLM results are reported with standard deviations. Human results are reported from Lauer and O’Brien (2020).
Experiment 2 evaluates whether MLLMs interpret consistently two visualizations that represent the same underlying data table, one misleading and the other non-misleading. This is assessed using a Likert-scale question, such as in Figure 2. Unlike in Experiment 1, there is no ground truth answer. Instead, since both visualizations depict the same underlying data, a model is considered misled if it assigns different scores to each of them. We assess the significance of the difference in ratings with a two-sided paired t-test (p-value (p) $\leq$ 0.05).
Figure 6 summarizes the average Likert-scale ratings from all 18 MLLMs for each visualization pair. Human results from prior work (Lauer and O’Brien, 2020) are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is statistically significant for the two visualization pairs involving truncated axis: one with bar charts (p = 5e-3) and one with line charts (p = 0.01). For the other two pairs, misrepresentation with bubble charts (p = 0.45) and 3D effects with pie charts (p = 0.55), the differences are not statistically significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, in particular for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s answers consistency.
3.3 Experiment 3 - Correction methods
Having established that MLLMs are vulnerable to misleading visualizations, we explore six correction methods aimed at improving their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline on misleading visualizations in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) (McNemar, 1947). Figure 7 shows the change in accuracy relative to Experiment 1 across the six methods: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization.
Figure 7: Change ($\Delta$) in accuracy (percentage points) compared to Experiment 1 using different inference-time correction methods. Statistically significant changes ($\leq$0.05) are hashed.
Figure 8: Table extraction for a non-misleading visualization and its misleading version with an inverted axis.
Two correction methods are effective: table-based QA and redrawing the visualization. The most effective approach by far to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Its strength lies in fully replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis, or misleading annotations. However, this method proves less effective against misleaders that persist even when the data is presented in table form, such as inappropriate item order and concealed uncertainty. Table-based QA comes with a notable cost for non-misleading visualizations. Two models experience significant decreases in accuracy, by up to 8.5 pp. This degradation is primarily due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost.
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations (Fan et al., 2022). It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib (Hunter, 2007), which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%), and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer and more detailed Python code. Unlike table-based QA, there are no significant decreases in accuracy on non-misleading visualizations, making it a more attractive option when preserving performance on non-misleading visualization is a priority.
Table extraction is a key intermediate step. Both table-based QA and redrawing the visualization depend heavily on the accuracy of an intermediate step: table extraction. Since the underlying data table is not available, it must be derived from the visualization before either correction method can be applied. This extraction step is often non-trivial, particularly for complex chart types such as maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. Perfect accuracy in table extraction is not always necessary, its importance depends on the nature of the question. Some questions can be answered correctly as long as the extracted table preserves the overall trend of the underlying data, while others require exact values. As discussed above, errors in table extraction can significantly decrease accuracy on non-misleading visualizations. In the absence of ground truth tables, we cannot directly evaluate the accuracy of table extraction. Instead, we analyze MLLM outputs on CHARTOM (Rho et al., 2023; Bharti et al., 2024), which pairs each question with two visualizations, one misleading and one non-misleading, generated from the same underlying data. Ideally, the tables extracted from both visualizations should be identical. However, for all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). In Figure 8, despite that the two visualizations represent the same data, the extracted tables show different values and trends. The table extracted from the inverted axis is incorrect. We make the hypothesis that misleaders negatively affect the quality of table extraction too. Given that the two most effective correction methods depend on reliable table extraction, and that extraction errors also negatively impact performance on non-misleading visualizations, improving this intermediate step is a key direction for future work.
Other correction methods are not effective. None of the remaining correction methods yield significant improvements. Including the extracted table in the prompt even results in a significant decrease in accuracy on non-misleading visualizations for one MLLM. While using the extracted table alone proves to be the most effective method, combining it with the original misleading visualization does not produce similar gains. This further suggests that removing or replacing the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method. Although many misleaders rely on axis manipulations, providing the extracted axes alone has little to no effect on performance. Similarly, adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task (Lo and Qu, 2025; Alexander et al., 2024). Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing to select the most suitable correction method based on the type of misleader.
4 Related work
Our correction methods, in particular redrawing the visualization, relate to prior work on visualization linters (Hopkins et al., 2020; Chen et al., 2022; Fan et al., 2022): rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations (Fan et al., 2022), they can only operate with specific chart design tools (Hunter, 2007; Hopkins et al., 2020; Chen et al., 2022) or a very limited set of chart and misleader types (Fan et al., 2022). In contrast, our correction methods do not impose such restrictions.
Our findings are further supported by prior work (Bendeck and Stasko, 2025) and parallel studies (Pandey and Ottley, 2025; Chen et al., 2025; Valentim et al., 2025), which also report MLLM vulnerabilities to misleaders (Bendeck and Stasko, 2025; Pandey and Ottley, 2025; Chen et al., 2025) and to other design decisions like the color palette (Valentim et al., 2025). Our study differentiates itself in several key ways. Prior work (Bendeck and Stasko, 2025) only evaluated GPT4 on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world visualizations and show that vulnerability of MLLMs is somewhat lower in these cases. Furthermore, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving QA performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works.
5 Conclusion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Experiment 1 demonstrated that MLLMs are unable to accurately interpret the underlying data when presented with misleading visualizations. Experiment 2 showed that MLLMs are particularly susceptible to truncated axes, often failing to provide consistent answers when shown misleading and non-misleading versions of the same data. Identifying this vulnerability fills a critical gap in automated chart understanding research. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation (Correll and Heer, 2017).
To mitigate this issue, we proposed six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, table-based QA comes at the cost of reduced accuracy on non-misleading visualizations. While most recent work in chart understanding treats visualizations purely as images (Zeng et al., 2025; Zhang et al., 2024; Masry et al., 2025), our findings challenge this direction and sheds a new light on alternative approaches that involve table extraction (Liu et al., 2023a, b; Kim et al., 2025).
6 Limitations
We identify three limitations to our work.
First, the visualization redrawing method does not support maps, as the Matplotlib library (Hunter, 2007) lacks sufficient functionality for rendering high-quality maps.
Second, we assume prior knowledge of the chart type (bar, line, …) for generating the prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step.
Third, existing datasets do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI (Ge et al., 2023) and present to a lower extent in the real-world cases (Lo et al., 2022).
Acknowledgments
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We want to express our gratitude to Niklas Traser for conducting an initial exploration of the real-world data, to Jan Zimny for our insightful discussions on the topic of misleading visualizations, and to Germàn Ortiz, Manisha Venkat, and Max Glockner for their feedback on a draft of this work.
Appendix A Definition of misleaders
Misleader Definition
Inverted axis ($n$=26) An axis is oriented in an unconventional direction
and the perception of the data is reversed (Lo et al., 2022).
Truncated axis ($n$=21) The axis does not start from zero or is truncated
in the middle resulting in an exaggerated difference
between the two bars (Lo et al., 2022).
Inappropriate axis range ($n$=15) The axis range is either too broad or too narrow to
accurately visualize the data, allowing changes to be
minimized or maximized depending on
the author’s intention (Lo et al., 2022).
Inconsistent tick intervals ($n$=12) Cases with varying intervals between the ticks (Lo et al., 2022).
3D effects ($n$=12) The closer something is, the larger it appears,
despite being the same size in 3D perspective (Lo et al., 2022).
Inappropriate item order ($n$=9) The axis labels or legends appear to be in a random
order due to manipulation of data ordering (Ge et al., 2023).
Inappropriate aggregation ($n$=8) Aggregating data in an improper way that leads to
inaccurate conclusions (Ge et al., 2023).
Dual axis ($n$=8) Two independent axes are layered on top of
each other with inappropriate scaling (Lo et al., 2022).
Misrepresentation ($n$=7) The value labels provided do not match
the visual encoding (Lo et al., 2022).
Cherry picking ($n$=6) Selecting only a subset of data to display, which
can be misleading if one is asked to infer something
about the whole set of data (Ge et al., 2023).
Misleading annotations ($n$=5) Annotations that contradict or make it harder
to read the visualization (Ge et al., 2023).
Area encoding ($n$=5) Linearly encoding the values as areas
leads the readers to consistently
underestimate the values (Lo et al., 2022).
Concealed uncertainty ($n$=3) Not displaying uncertainty in visualizations may
misrepresent the certainty in the underlying data.
In the case of prediction making, this can misguide
the viewers to falsely overconfident conclusions (Ge et al., 2023).
Missing normalization ($n$=3) Displaying unnormalized data in absolute quantity
when normalized data in relative quantity
is of interest (Ge et al., 2023).
Inappropriate use of pie chart ($n$=1) When a pie chart is used for non-part-to-whole data,
it creates confusion for the audience, who may
misinterpret the significance of a given section (Lo et al., 2022).
Missing data ($n$=1) A visual representation implies data exist but
the data is actually missing (Ge et al., 2023).
Overplotting ($n$=1) Displaying too many things on a plot can
obscure parts of the data (Ge et al., 2023).
Table 2: The misleaders included in this work, with their number of occurrences ($n$).