### v5

# Protecting multimodal large language models against misleading visualizations

Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, i.e., charts that distort the underlying data, leading readers to draw inaccurate conclusions. Here, we uncover an important vulnerability: MLLM question-answering (QA) accuracy on misleading visualizations drops on average to the level of the random baseline. To address this, we provide the first comparison of six inference-time methods to improve QA performance on misleading visualizations, without compromising accuracy on non-misleading ones. We find that two methods, table-based QA and redrawing the visualization, are effective, with improvements of up to 19.6 percentage points. We make our code and data available.111github.com/UKPLab/arxiv2025-misleading-visualizations
Protecting multimodal large language models against
misleading visualizations
Jonathan Tonglet1,2,3, Tinne Tuytelaars2, Marie-Francine3 Moens Iryna Gurevych1 1 Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science, TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE 2 Department of Electrical Engineering, KU Leuven 3 Department of Computer Science, KU Leuven www.ukp.tu-darmstadt.de
1 Introduction
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, and companies to efficiently communicate data insights to a broad audience (Huang et al., 2025). They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public (Zhang et al., 2021). In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers (Pandey et al., 2014). However, visualizations can also be misleading (Correll and Heer, 2017). This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Correll and Heer, 2017; Lauer and O’Brien, 2020; McNutt et al., 2020; Yang et al., 2021; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). These misleaders appear across a wide range of visualizations, including bar and line charts, or choropleth maps. Recent taxonomies have documented over 70 types of misleaders observed in real-world examples (Lo et al., 2022; Lan and Liu, 2025). Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like 3D effects. Figure 1 shows three real-world examples of misleading visualizations (Lo et al., 2022) paired with multiple choice questions (MCQs). Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer instead of the correct one. Figure 2 illustrates how two visualizations of the same data, one non-misleading and the other misleading, can lead to different interpretations by readers.
Figure 1: Three examples of real-world misleading visualizations (Lo et al., 2022) with MCQs. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
Figure 2: Non-misleading and misleading visualizations of the same data (Lauer and O’Brien, 2020) with a Likert-scale question where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses. However, the deceived reader chooses a higher value for the truncated bar chart.
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies (Pandey et al., 2015; Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024) have shown that such visualizations reduce human accuracy when answering MCQs about the underlying data. Other studies (Pandey et al., 2015; O’Brien and Lauer, 2018; Lauer and O’Brien, 2020) have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
Misleading visualizations pose a serious threat to society. Due to their deceptive potential, they can be exploited by malicious actors to promote online disinformation (Correll and Heer, 2017). For instance, the charts in Figure 1 mislead readers on sensitive topics such as COVID-19, abortion, and gun violence. Misleading visualizations can shift public opinion, even on polarized debates like Brexit (Tartaglione and de Wit, 2025).
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research (Huang et al., 2025), with steady progress on reference benchmarks such as ChartQA (Masry et al., 2022; Hoque et al., 2022). However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it. These risks underscore the urgency of thoroughly assessing the vulnerability of MLLMs to misleading visualizations and developing effective mitigation strategies.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 19 MLLMs of varying sizes to misleading visualizations. In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs on misleading and non-misleading visualizations. In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. The results of both experiments demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. In Experiment 3, we compare six inference-time correction methods to reduce the vulnerability of MLLMs to the misleading visualizations of Experiment 1.
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 19 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We propose the first analysis of six correction methods to mitigate the negative impact of misleaders on QA tasks.
2 Methodology
2.1 Datasets
Experiments 1 and 3 We compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Appendix A (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lo et al., 2022); (b) a non-misleading visualization dataset ($n$ = 124) (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lee et al., 2017), also combining real-world and synthetic cases; and (c) ChartQA (Masry et al., 2022), the reference real-world non-misleading benchmark ($n$ = 2500). Datasets (a) and (b) combine three existing resources and one introduced in this work. First, CALVI (Ge et al., 2023) includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with an MCQ. CALVI was initially designed and is the reference resource for evaluating humans. Second, CHARTOM (Rho et al., 2023; Bharti et al., 2024) contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. Third, VLAT (Lee et al., 2017), the reference dataset to evaluate humans on non-misleading cases, provides 12 visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with an MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them (Lo et al., 2022), which inspired the synthetic examples in CALVI. We manually create MCQs, using as a reference those from CALVI and CHARTOM. By incorporating visualizations and questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
Experiment 2 We use four pairs of visualizations, one misleading and the other non-misleading, designed by Lauer and O’Brien (2020).
Table 1 provides detailed dataset statistics.
Dataset Experiments $n$ $m$ Question types
Misleading visualizations 1 & 3 143 17 MCQ, Free-text, rank
$\hookrightarrow$ CALVI (misleading) 1 & 3 45 14 MCQ
$\hookrightarrow$ CHARTOM (misleading) 1 & 3 56 7 MCQ, Free-text, rank
$\hookrightarrow$ Real-world 1 & 3 42 12 MCQ
Non-misleading visualizations 1 & 3 124 - MCQ, Free-text, rank
$\hookrightarrow$ CALVI (non-misleading) 1 & 3 15 - MCQ
$\hookrightarrow$ CHARTOM (non-misleading) 1 & 3 56 - MCQ, Free-text, rank
$\hookrightarrow$ VLAT 1 & 3 53 - MCQ
ChartQA 1 2500 - Free-text
Lauer & O’Brien 2 8 3 Likert-scale
Table 1: Datasets statistics. $n$ is the number of instances. $m$ is the number of misleader categories.
2.2 Correction methods
Figure 3: Illustration of the six inference-time correction methods applied to a misleading visualization from CALVI (Ge et al., 2023). The visualization suffers from inconsistent tick intervals on the y-axis.
In Experiment 3, we compare six inference-time correction methods, illustrated in Figure 3, to increase QA accuracy on misleading visualization without compromising it on non-misleading ones.
(1) Misleader warning: we insert in the prompt a message to make the MLLM aware of the misleader in the visualization. The message is identical for all instances with the same misleader.
Providing (2) the axes data, (3) the data table, (4) or both: we prompt the MLLM in zero-shot to extract the axes labels or underlying data table. The axes and tables are formatted as text strings. We do not impose formatting constraints on the axes and tables. The axes, the table, or both are provided as additional prompt input.
(5) Table-based QA: after extracting the table with the MLLM, we provide it alone to a text-only LLM, reframing the task as table-based QA (Liu et al., 2023a, b; Kim et al., 2025).
(6) Redrawn visualization: after extracting the table with the MLLM, we provide it to a text-only LLM, which generates Python code to create a visualization using Matplotlib (Hunter, 2007). If the code compiles successfully, the generated visualization replaces the original one in the QA prompt; otherwise, the original one is used.
2.3 Implementation details
We conduct experiments with 19 MLLMs on a machine with two A100 GPUs, including 11 open-weight MLLMs from the Llava-Next (Liu et al., 2024), Qwen2VL (Wang et al., 2024), Ovis1.6 (Lu et al., 2024), and InternVL2.5 (Chen et al., 2024) families. Additionally, we include five commercial models, GPT4 and GPT4o (OpenAI, 2023), Gemini-1.5-flash and -pro (Gemini-Team, 2024), and Claude-3.5-sonnet (Anthropic, 2024), as well as three open-weight MLLMs specialized in chart understanding: Llava-Chart-Instruct (Zeng et al., 2025), TinyChart (Zhang et al., 2024), and ChartGemma (Masry et al., 2025). Qwen2.5-7B (Qwen-Team, 2024) serves as the LLM for table-based QA and visualization redrawing. We use the transformers library (Wolf et al., 2020) to access all open-weight models. All prompts are provided in Appendix B. We set the temperature to 0. Following the standard ChartQA evaluation setup, all (M)LLMs are prompted in zero-shot. We use the Direct approach for TinyChart (Zhang et al., 2024).
3 Results
3.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Figure 4: Top: Accuracy (%) of various MLLMs on misleading visualization, non-misleading visualization, and ChartQA datasets. The horizontal dashed line indicates the accuracy of the random baseline on misleading visualizations. Models are sorted by increasing accuracy on ChartQA. Bottom: Accuracy (%) of various MLLMs on subsets of the misleading visualizations. The horizontal dashed lines indicate average human accuracy on CALVI and CHARTOM (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024).
Experiment 1 assesses the ability of MLLMs to answer questions about the underlying data given a visualization. The upper part of Figure 4 presents the QA accuracy of 19 MLLMs across the three datasets, sorted by increasing accuracy on ChartQA. The bottom part displays the accuracy for the three subsets constituting the misleading visualizations dataset. The results of Experiment 1 reveal three key findings.
Performance is worse on misleading visualizations. All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 percentage points (pp), with an average of 23.8 pp. The decline is even more pronounced compared to ChartQA, reaching up to 65.5 pp, with an average of 53.8 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (26.4%) is comparable to the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret data distorted by misleaders.
For comparison, average human accuracy reported on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024). On the same datasets, MLLMs achieve an average of 52 and 48% on non-misleading visualizations and 24 and 22% on misleading ones, falling short of human performance in all cases.
The ranking of MLLMs by accuracy varies a lot across the three datasets. The two best-performing models on the CALVI subset of misleading visualizations rank among the worst on ChartQA, placing second and third from the bottom. This suggests that achieving high accuracy on misleading visualizations is unlikely to emerge naturally from improvements on ChartQA, underscoring the need for dedicated mitigation methods.
Performance is higher on real-world misleading visualizations than on other subsets. MLLMs perform better on the real-world subset, achieving an average accuracy of 34.7%, compared to 24% on CALVI and 22% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, Claude-3.5-sonnet, GPT4o, and Gemini-1.5-pro, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events, without relying on the content of the visualization. Appendix C discusses further the impact of parametric knowledge.
Performance varies by type of misleader. Not all misleading visualizations are equally deceptive. This was already observed with human readers (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024), and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Appendix D. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 63.2, 50.5, and 44.4, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 25.1 and 28.1 %, respectively. There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader (Rho et al., 2023). Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects (Ge et al., 2023). The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM (Rho et al., 2023), compared to MLLMs, which achieve an average of 35.5%. These results highlight a notable difference in how humans and MLLMs are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers (Fan et al., 2022).
3.2 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Experiment 2 evaluates whether MLLMs interpret consistently two visualizations that represent the same underlying data table, one misleading and the other non-misleading. This is assessed using a Likert-scale question, such as in Figure 2. Unlike in Experiment 1, there is no ground truth answer. Instead, since both visualizations depict the same underlying data, a model is considered misled if it assigns different scores to each of them. We assess the significance of the difference in ratings with a one-sided paired t-test (p-value (p) $\leq$ 0.05).
Figure 5 summarizes the average Likert-scale ratings from all 19 MLLMs for each visualization pair. Average human results from prior work (Lauer and O’Brien, 2020) are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is significant for the two visualization pairs involving truncated axis: one with bar charts (p = 3e-3) and one with line charts (p = 4e-3). For the other two pairs, misrepresentation with bubble charts (p = 0.23) and 3D effects with pie charts (p = 0.27), the differences are not significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, particularly for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s consistency.
Figure 5: Average Likert-scale ratings (1 to 6) on four pairs of misleading and non-misleading visualizations depicting the same data. MLLM results are reported with standard deviations. Average human results are reported from Lauer and O’Brien (2020).
3.3 Experiment 3 - Correction methods
[…]
Having established that MLLMs are vulnerable to misleading visualizations, we examine six correction methods designed to enhance their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) (McNemar, 1947). Figure 6 shows the change in accuracy relative to Experiment 1 across the six methods: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization.
Two correction methods are effective: table-based QA and redrawing the visualization. The most effective approach to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Originally proposed as a general approach for chart understanding (Liu et al., 2023a, b), table-based QA further shows its effectiveness for countering misleading charts. Its strength lies in replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis. However, this method proves less effective against misleaders that persist even when the data is presented in tabular format, such as inappropriate item order. Table-based QA comes with a significant cost for two models on non-misleading visualizations, by up to 8.5 pp. This degradation is due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost. Appendix E reports table-based QA performance using DePlot (Liu et al., 2023a) and MatCha (Liu et al., 2023b), two methods developed specifically for table extraction.
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations (Fan et al., 2022). It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib (Hunter, 2007), which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%) and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer Python code. Unlike table-based QA, there are no significant decreases in accuracy on non-misleading visualizations, making it a more attractive option when preserving performance on non-misleading visualizations is a priority.
Appendix F discusses the impact of combining these two effective methods with Chain-of-Thought prompting (Wei et al., 2022).
Table extraction is a key intermediate step. Both table-based QA and redrawing the visualization depend heavily on the accuracy of the intermediate table extraction step. This extraction step is often non-trivial, particularly for maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. This is particularly prevalent for inverted axis, 3D effects, and dual axis, as illustrated in Appendix G. Perfect accuracy in table extraction is not always necessary; its importance depends on the nature of the question. Some questions can be answered correctly if the extracted table preserves the overall trend of the underlying data, while others require exact values. We conduct a comprehensive qualitative analysis of table extraction in Appendix H.
Other correction methods are not effective. None of the remaining correction methods yields significant improvements. While using the extracted table alone proves to be the most effective method, combining it with the misleading visualization does not produce similar gains. The MLLMs are biased towards using the image and ignoring the table in most cases.
Although many misleaders rely on axes manipulations, providing the extracted axes has no significant effect on performance. A manual error analysis showed that the axes extraction step is more often accurate than table extraction, as reported in Appendix I. However, unlike the table, the axes alone are insufficient to answer the question. The MLLM still needs to combine the visualization image with the cues from the axes that indicate the presence of misleaders. The results show that this combination is challenging. This further suggests that removing or modifying the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method.
Adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task (Lo and Qu, 2025; Alexander et al., 2024). Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing the selection of the most suitable correction method based on the type of misleader.
4 Related work
The correction methods, particularly misleader warnings and redrawing the visualization, relate to prior work on visualization linters (Hopkins et al., 2020; Chen et al., 2022; Fan et al., 2022): rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations (Fan et al., 2022), they can only operate with specific chart design tools (Hunter, 2007; Hopkins et al., 2020; Chen et al., 2022) or a very limited set of chart and misleader types (Fan et al., 2022). In contrast, the correction methods we explored do not impose such restrictions. Table-based QA was originally introduced as a general approach for chart understanding by Liu et al. (2023a, b). Although this approach has been surpassed on standard benchmarks by more recent methods that treat visualizations purely as images Zeng et al. (2025); Masry et al. (2025), our results reveal that table-based QA finds a new purpose as an effective method for counteracting misleading visualizations.
Our findings are further supported by prior work (Bendeck and Stasko, 2025) and parallel studies (Pandey and Ottley, 2025; Chen et al., 2025; Valentim et al., 2025), which also report MLLM vulnerabilities to misleaders (Bendeck and Stasko, 2025; Pandey and Ottley, 2025; Chen et al., 2025) and to other design decisions like the color palette (Valentim et al., 2025). Our study differentiates itself in several key ways. Prior work (Bendeck and Stasko, 2025) only evaluated GPT4 on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world visualizations and show that the vulnerability of MLLMs is lower in these cases. Furthermore, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving QA performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works.
5 Conclusion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Identifying this vulnerability fills a critical gap in the research on automated chart understanding. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation (Correll and Heer, 2017). To mitigate this issue, we evaluated six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, correction methods often come at the cost of reduced accuracy on non-misleading visualizations.
Limitations
We identify three limitations to our work.
First, the visualization redrawing method does not support maps, as the Matplotlib library (Hunter, 2007) lacks sufficient functionality for rendering high-quality maps.
Second, we assume prior knowledge of the chart type (e.g., bar, line) for generating prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step.
Third, existing misleading visualization datasets are small compared to standard datasets in the automated chart understanding literature, such as ChartQA. Furthermore, they do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI (Ge et al., 2023) and, to a lesser extent, in real-world data. Despite these limitations, CALVI is the reference dataset for assessing human vulnerability to misleading visualizations. It was carefully curated by experts in data visualization and validated on a large sample of human subjects. Therefore, it constitutes a valid resource for evaluating MLLMs. The properties of CALVI are similar to those of VLAT (Lee et al., 2017), an expert-created dataset for non-misleading chart understanding, which acts as a reference despite being much smaller than datasets such as ChartQA. Furthermore, the real-world distribution of misleaders is inherently imbalanced (Lo et al., 2022), with categories such as truncated axis being more prevalent than others. Hence, scaling the real-world dataset would not address the imbalance issues of synthetic datasets, such as CALVI.
Ethics statement
Social impact
Misleading visualizations are a prevalent form of multimodal misinformation. Our results show the importance of protecting not only humans but also MLLMs from misleading visualizations. Our correction methods provides the first promising results in that direction.
Risks
While the proposed correction methods improve the performance of MLLMs on misleading visualizations, they do not guarantee a correct answer. In general, human users should not blindly trust the output of chart understanding systems.
Dataset access
We release the CALVI, VLAT, Likert-scale, and real-world datasets under a CC-BY-SA-4.0 license. The real-world dataset is only intended to be used for research purposes. CHARTOM can be accessed by contacting its authors (Bharti et al., 2024).
AI assistants use
AI assistants were used in this work to assist with writing by correcting grammar mistakes and typos.
Acknowledgments
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We want to express our gratitude to Niklas Traser for conducting an initial exploration of the real-world data, to Jan Zimny for our insightful discussions on the topic of misleading visualizations, and to Germàn Ortiz, Manisha Venkat, and Max Glockner for their feedback on a draft of this work.
Appendix A Definition of misleaders
Table 2 provides the definitions of all the misleaders considered in this work (Lo et al., 2022; Ge et al., 2023), with their number of occurences.
Misleader Definition
Inverted axis (nn=26) An axis is oriented in an unconventional direction
and the perception of the data is reversed (Lo et al., 2022).
Truncated axis (nn=21) The axis does not start from zero or is truncated
in the middle resulting in an exaggerated difference
between the two bars (Lo et al., 2022).
Inappropriate axis range (nn=15) The axis range is either too broad or too narrow to
accurately visualize the data, allowing changes to be
minimized or maximized depending on
the author’s intention (Lo et al., 2022).
Inconsistent tick intervals (nn=12) Cases with varying intervals between the ticks (Lo et al., 2022).
3D effects (nn=12) The closer something is, the larger it appears,
despite being the same size in 3D perspective (Lo et al., 2022).
Inappropriate item order (nn=9) The axis labels or legends appear to be in a random
order due to manipulation of data ordering (Ge et al., 2023).
Inappropriate aggregation (nn=8) Aggregating data in an improper way that leads to
inaccurate conclusions (Ge et al., 2023).
Dual axis (nn=8) Two independent axes are layered on top of
each other with inappropriate scaling (Lo et al., 2022).
Misrepresentation (nn=7) The value labels provided do not match
the visual encoding (Lo et al., 2022).
Cherry picking (nn=6) Selecting only a subset of data to display, which
can be misleading if one is asked to infer something
about the whole set of data (Ge et al., 2023).
Misleading annotations (nn=5) Annotations that contradict or make it harder
to read the visualization (Ge et al., 2023).
Area encoding (nn=5) Linearly encoding the values as areas
leads the readers to consistently
underestimate the values (Lo et al., 2022).
Concealed uncertainty (nn=3) Not displaying uncertainty in visualizations may
misrepresent the certainty in the underlying data.
In the case of prediction making, this can misguide
the viewers to falsely overconfident conclusions (Ge et al., 2023).
Missing normalization (nn=3) Displaying unnormalized data in absolute quantity
when normalized data in relative quantity
is of interest (Ge et al., 2023).
Inappropriate use of pie chart (nn=1) When a pie chart is used for non-part-to-whole data,
it creates confusion for the audience, who may
misinterpret the significance of a given section (Lo et al., 2022).
Missing data (nn=1) A visual representation implies data exist but
the data is actually missing (Ge et al., 2023).
Overplotting (nn=1) Displaying too many things on a plot can
obscure parts of the data (Ge et al., 2023).
Table 2: The misleaders included in this work, with their number of occurrences ($n$).
Appendix B Prompts
Figures 8, 7, and 9 provide the prompt for the intermediate metadata extraction and visualization redrawing steps. Figure 10 provides the QA prompt variants used in Experiments 1, 2, and 3. Parts of the prompt are included only for certain question types and correction methods.
Table extraction prompt Generate the underlying data table of the figure below. Change columns with |, change row by starting a new line. Provide only the table as output.
Figure 7: Table extraction prompt.
Axes extraction prompt For maps: What is the legend and its categories (with their colors) in this map? Answer only with the content of the legend and its categories (with their colors). For pie charts: What are the categories in this pie chart? Answer only with the categories. Other chart types: What are the axis labels and ticks of this chart? Answer only with the axis labels and the tick values, going from the bottom-left to the top-left corner of the chart for the y-axis and from the bottom-left to the bottom-right corner for the x-axis.
Figure 8: Axes extraction prompt.
Visualization redrawing prompt Generate the matplotlib code to generate this {CHART_TYPE}, using the tabular data below. Provide only the code as output, including the table values represented as list or numpy arrays. {TABLE}
Figure 9: Visualization redrawing prompt.
QA prompt If table-based QA: {TABLE} {QUESTION} If MCQ: Provide the correct answer among the following choices: {CHOICES} If rank: Provide the answer as a Python list. If misleader warning: Be careful, the following design flaw has been identified in the chart: {MISLEADER DEFINITION} If axes: Below is a description of the charts axis labels or legend. {AXIS} If table: Below is a table containing the values represented in the chart. {TABLE} Provide only the final answer to the question.
Figure 10: QA prompt.
Figure 11: Change ($\Delta$) in accuracy (percentage points) compared to Experiment 1 using Chain-of-Thought (CoT) prompting alone or in combination with correction methods.
Figure 12: QA Accuracy (%) per type of misleader.
Appendix C Experiment 1 - Impact of parametric knowledge
We observed in Experiment 1 that Claude-3.5-sonnet, GPT4o, and Gemini-1.5-pro primarily outperform other MLLMs on the real-world subset of the misleading visualizations. We make the assumption that this is due to their parametric knowledge, which allows them to answer the question without considering the visualization. To support this assumption, we identified the subset of 22 real-world MCQS, out of 42, that could be answered based on world knowledge. We rephrased the questions slightly to make them self-contained, without direct references to the visualization’s content. The accuracy of GPT4o on this subset is 77%. If we provide only the MCQ to GPT4o, without the visualization’s image, the accuracy is still 50%, highlighting indeed a moderate ability to answer the real-world MCQs based on parametric knowledge. In such cases, parametric knowledge effectively serves as a form of protection against real-world misleaders. However, this finding should be interpreted with caution. In practical scenarios, the underlying data will often be recent and unlikely to be covered by the parametric knowledge.
Appendix D Experiment 1 - accuracy by misleader
Figure 12 provides the accuracy per misleader for each MLLM in Experiment 1.
Appendix E Experiment 3 - Table-based QA results with DePlot and Matcha
Model Misleading visualizations Non-misleading visualizations
DePlot 44.1 46.0
MatCha 42.7 37.9
Qwen2VL-7B 42.0 41.1
Ovis1.6-9B 39.9 47.6
InternVL2.5-8B 44.1 52.4
Table 3: Table-based QA accuracy using Qwen2.5-7B, with different models for table extraction (%).
Table 3 compares the performance of table-based QA using MLLMs with that of two smaller specialized chart-to-table extraction models, DePlot (Liu et al., 2023a) and MatCha (Liu et al., 2023b). On misleading visualizations, the specialized models outperform Ovis1.6-9B and Qwen2VL-7B. Only InternVL2.5-8B matches the performance obtained with DePlot. This indicates that smaller models fine-tuned specifically for table extraction constitute a strong alternative to MLLMs for countering misleading visualizations.
However, on non-misleading visualizations, table-based QA performs substantially worse with DePlot compared to InternVL2.5-8B. Moreover, MatCha achieves the lowest performance among all models. We attribute this gap to the high diversity of chart types in the non-misleading dataset, particularly in the VLAT subset, which includes complex visualizations such as bubble charts and treemaps. Since DePlot and MatCha were not fine-tuned on such chart types, they lack the broad generalization capabilities of larger MLLMs, which explains their weaker performance in this setting.
Figure 13: Table extraction for a non-misleading visualization and its misleading version with an inverted axis.
Figure 14: Table extraction for a non-misleading visualization and its misleading version with 3D effects.
Appendix F Experiment 3 - Results with Chain-of-Thought prompting
By default, we evaluate MLLMs in a direct prompting setting where they only output the final answer. Prompting techniques such as Chain-of-Thought (CoT), which requires the MLLM to generate intermediate reasoning steps before providing the final answer, are known to improve performance on several reasoning tasks (Wei et al., 2022).
Figure 11 shows the change in accuracy when using zero-shot CoT prompting, either alone or in combination with one of the two most effective correction methods, table-based QA and redrawing the visualization. CoT alone improves performance on both misleading and non-misleading visualizations. The gains are generally modest, except for InternVL2.5-8B, where the improvement exceeds 10 pp on misleading visualizations.
Using CoT has no significant impact on misleading visualizations for table-based QA. However, it substantially reduces the negative impact of table-based QA on non-misleading visualizations, bringing the performance drop close to 0 for Qwen2VL-7B and even bringing a net positive change in accuracy for InternVL2.5-8B.
Combining CoT with redrawn visualizations worsens accuracy on non-misleading visualizations. Its effect on misleading visualizations is model-dependent. Ovis1.6-9B shows a small decrease, while InternVL2.5-8B experiences a large performance gain.
Overall, CoT alone is not an effective correction method for misleading visualizations. It can be beneficial when combined with table-based QA, mainly by mitigating the negative impact on non-misleading data. However, its effects are inconsistent across models and settings, and it should therefore be applied with caution.t be combined with redrawing the visualization.
Figure 15: Table extraction for a non-misleading visualization and its misleading version with dual axis.
Appendix G Experiment 3 - Examples of errors in table extraction due to misleaders
Figures 13, 14, and 15 illustrate representative examples of table extraction errors made by a MLLM due to misleaders. The examples are taken from CHARTOM, where each data table is associated with two visualizations, one misleading and one non-misleading. In all three cases, the table extracted from the non-misleading visualization is accurate. However, the MLLM fails to recover the same table from the misleading version.
In Figure 13, the MLLM fails to correctly align the dots with the inverted axis. In Figure 14, the 3D effects distort the perceived size of the bars on the right side, leading to incorrect value extraction. In Figure 15, the MLLM first extracts values for both lines using the left axis, and then switches to the right axis for 2022, failing to recognize that each line is associated with a different axis.
Appendix H Experiment 3 - Analysis of table extraction
Impact of table extraction on table-based QA
We conduct a manual analysis on a random sample of 30 misleading visualizations to assess the impact of the table extraction step on the accuracy of table-based QA. We analyze the three MLLMs from Experiment 3 as well as DePlot, introduced in Appendix E. The results are reported in Table 4. In most cases, both the table extraction and the QA step are incorrect. For all models except DePlot, the second most frequent outcome is that both steps are correct.
In some instances, the extracted table is incorrect but the QA output is correct. This typically occurs when the correct answer is “not enough information” and is selected not because of a correct reasoning but because there are missing or incorrect values in the extracted table.
In other cases, the table extraction is correct but the QA is incorrect. Except for Ovis1.6-9B, this category is the least prevalent. This occurs mostly with misleaders that are not removed by converting the visualization to the table modality. For example, if the misleader is inappropriate item order, the entries in the table remain shuffled, which is deceiving. If the misleader is cherry-picking, there is still not enough data to answer the question properly.
Table-based QA (Qwen2.5-7B)
Table extraction Correct Incorrect
Qwen2VL-7B Correct 23.3 10.0
Incorrect 16.7 50.0
Ovis1.6-9B Correct 26.7 20.0
Incorrect 20.0 33.3
InternVL2.5-8B Correct 33.3 20.0
Incorrect 10 36.7
DePlot Correct 13.3 10.0
Incorrect 20.0 56.7
Table 4: Manual analysis of the impact of table extraction accuracy on table-based QA accuracy, on a random sample of 30 misleading visualizations (%).
Table extraction accuracy by misleader
Table 5 reports our manual analysis of table extraction accuracy with InternVL2.5-8B, the strongest MLLM for table-based QA. We analyze all instances of the ten most frequent misleaders in the misleading visualization dataset. The results show that extraction accuracy varies substantially across misleaders.
For some misleaders, such as truncated axis and misrepresentation, the extraction accuracy exceeds 80%. The distortions introduced by these misleaders have a limited impact on table extraction. For truncated axis, it is sufficient to align the top of the bars with the y-axis to recover the values. For misrepresentation, the model often succeeds because the relevant value is explicitly displayed on the bar or pie slice.
In contrast, several misleaders have a severe negative impact on table extraction. When a dual axis is present, all extracted tables are incorrect, as illustrated in Figure 15. Inverted axis and 3D effects also lead to very low extraction accuracy.
Overall, these results show that table-based QA is not a universal correction method for all misleaders. Future work should either improve table extraction for challenging misleaders such as inverted and dual axis, for example through creating large scale synthetic data and fine-tuning specialized models like DePlot, or design dedicated correction methods for these misleaders.
Misleader Table extraction accuracy
Inverted axis 15.4
Truncated axis 81.0
Inappropriate axis range 46.7
Inconsistent tick intervals 50.0
3D 8.3
Inappropriate item order 33.3
Inappropriate aggregation 37.5
Dual axis 0.0
Misrepresentation 85.7
Cherry-picking 66.7
Table 5: Manual analysis of table extraction accuracy per misleader category, using Internvl2.5-8B (%).
Consistence of table extraction on CHARTOM
We conducted a manual analysis of the table extraction outputs of all three MLLMs on CHARTOM, which pairs each question with two visualizations, one misleading and one non-misleading, generated from the same underlying data. Ideally, the tables extracted from both visualizations should be identical. However, for all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). This further shows the negative impact of misleaders on table extraction accuracy, as illustrated in the examples of Appendix G.
Appendix I Experiment 3 - Analysis of axes extraction
Visualization + axes QA
Table extraction Correct Incorrect
Qwen2VL-7B Correct 16.7 50.0
Incorrect 13.3 20.0
Ovis1.6-9B Correct 20.0 56.7
Incorrect 10.0 13.3
InternVL2.5-8B Correct 23.3 60.0
Incorrect 6.7 10.0
Table 6: Manual analysis of the impact of axes extraction accuracy on visualization+axes QA accuracy, on a random sample of 30 misleading visualizations (%).
We conduct a manual analysis on a random sample of 30 misleading visualizations to assess the impact of the axes extraction step on the accuracy of the visualization+axes correction method. The results are reported in Table 6.
Compared to the table extraction results shown in Table 4, axes extraction is often accurate, with 20 to 25 out of 30 instances containing correct axis information. However, this high extraction accuracy does not translate into strong QA performance. For all MLLMs, at least half of the instances exhibit correct axes but incorrect QA predictions. Two main factors explain this behavior. First, the MLLMs often appear to ignore the extracted axes information and remain primarily influenced by the image modality. Second, the axes are unrelated to the distortions introduced by several misleaders, such as 3D effects or misrepresentation, which limits their corrective value.
Axes extraction errors occur most frequently for inconsistent tick intervals, where the MLLMs tend to hallucinate evenly spaced ticks along the axis.